Alpha particles (charge = +2e, mass = 6.68 x 10^-2
Answer #1

F = q E = 1.6 * 10-19 * 20800 = 3.33 * 10-15 N

FE in -y direction , FE = -3.33 * 10-15 N

FB = q v B sin theta

= 1.6 * 10-19 * 2.6 * 104 * 0.80 * sin 90

= 3.33 * 10-15 N

Ftotal = FB + FE

= 3.33 * 10-15 N + [- 3.33 * 10-15 N]

= -7.0 * 10-15 N